OpenStudy (anonymous):

How would I solve this: sqrt(-2 x)-(sqrt(5-x)) = -3/(sqrt(5-x))

OpenStudy (anonymous):

i would start by multiplying both sides by \(\sqrt{5-x}\) to get \[\sqrt{2x(5-x)}-(5-x)=-3\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I've done that

OpenStudy (anonymous):

now rewrite as \[\sqrt{-2x(5-x)}=-3+5-x=2-x\]

OpenStudy (anonymous):

square both sides and get \[-2x(5-x)=(2-x)^2\]

OpenStudy (anonymous):

then \[2x^2-10x=x^2-4x+4\] and you have a quadratic equation to solve

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

\[x^2-6x-4=0\]

OpenStudy (anonymous):

x=6.60555127546399 x=−0.6055512754639891

OpenStudy (anonymous):

then easiest to complete the square \[x^2-6x=4\] \[(x-3)^2=4+9=13\] \[x-3=\pm\sqrt{13}\] \[x=3\pm\sqrt{13}\]

OpenStudy (anonymous):

you can check that this is right here https://www.wolframalpha.com/input/?i=sqrt%28-2+x%29-%28sqrt%285-x%29%29+%3D+-3%2F%28sqrt%285-x%29%29

OpenStudy (anonymous):

cool, thanks so much!